OpenStudy (anonymous):

A sequence {An}is defined A(1)=0&A(n+1)=A(n)+2(n)WHAT IS VALUE OF A(100) (a)9902(B)9900(C)10100(D)9904

OpenStudy (anonymous):

In one site i got solution... but i did not understant..... Last will be A (1) +2*99*98*.....1 Now put the value 1+2(1+2+......+99) 2+ 2*99(99+1)/2 =9902

OpenStudy (anonymous):

plz any one understand explain me...

OpenStudy (anonymous):

a[1] = 0; a[2] = 2 (1) a[3] = 2 (1) + 2 (2) a[4] = 2 (1) + 2 (2) + 2 (3) a[n + 1] = 2 ( 1 + 2 + n) = n (n + 1) a[100] = 99 x100=9900

OpenStudy (anonymous):

No the answer is 9902...

OpenStudy (anonymous):

Is a(1)=0

OpenStudy (anonymous):

ya...

OpenStudy (anonymous):

So what is wrong with my computation above?

OpenStudy (anonymous):

In my book given answer is 9902...

OpenStudy (anonymous):

they have given the solution.... but i did not understand it.... Last will be A (1) +2*99*98*.....1 Now put the value 1+2(1+2+......+99) 2+ 2*99(99+1)/2 =9902

OpenStudy (anonymous):

Either the problem is not stated right or the answer in your book is wrong

OpenStudy (anonymous):

kk thankyou....:-)

OpenStudy (anonymous):

Do you know the the formula \[ 1+1 +3 + \cdots + n= \frac{n(n+1)}2 \]

OpenStudy (anonymous):

ya i know....

OpenStudy (anonymous):

\[ 1+2 +3 + \cdots + n= \frac{n(n+1)}2 \]

OpenStudy (anonymous):

ya...

OpenStudy (anonymous):

That is all what you need to do your problem

OpenStudy (anonymous):

kk... i think my solution is wrong.... thanx for your help...